OpenStudy (76yz):

solve the inequality, please explain too 2|4x+1|-5<1

OpenStudy (adilalvi):

2|4x+1|-5<1 8x+2-5<1 8x-3<1 8x<1+3 8x<4 now divide both side by 8 x=0.5 answer

OpenStudy (adilalvi):

thats what i know how to solve it is it right?

OpenStudy (76yz):

Looks correct to me, also quick question, could I instead of adding 1+3, do this: 8x-3<1 +3 +3 8x=4 =0.5

OpenStudy (adilalvi):

yes correct.

OpenStudy (76yz):

both would always give the right answer?

OpenStudy (adilalvi):

usually it means that what you do on one side you have to do it on another side.

OpenStudy (kevin):

2|4x+1|-5<1 |8x + 2| < 6 ---------------------------------------- 8x + 2 = 0 8x = -2 x = -1/4 for x >= 0 -(8x + 2) = 0 -8x = 2 x = -1/4 for x < 0 ----------------------------------------- 2|4x+1|-5<1 |8x + 2| < 6 i (8x + 2) < 6 8x < 4 x < 1/2 ii -(8x + 2) < 6 -8x - 2 < 6 -8x < 8 -x < 1 x > -1 So the answer is -1 < x < 1/2